Consider the set B={(1,1,1),(0,2,2),(0,0,3)}.show that B

I) spans R3

ii)is linearly independent

iii)is a basis for R3

1
Expert's answer
2021-08-15T18:16:29-0400

i)


"\begin{pmatrix}
 1 & 0 & 0 \\
 1 & 2 & 0 \\
 1 & 2 & 3 \\
\end{pmatrix}"

"R_2=R_2-R_1"


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 2 & 0 \\
 1 & 2 & 3 \\
\end{pmatrix}"

"R_3=R_3-R_1"


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 2 & 0 \\
 0 & 2 & 3 \\
\end{pmatrix}"

"R_2=R_2\/2"


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 2 & 3 \\
\end{pmatrix}"

"R_3=R_3-2R_2"


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 3 \\
\end{pmatrix}"

"R_3=R_3\/3"


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 1 \\
\end{pmatrix}"

The rank of the matrix is 3, so the given vectors span a subspace of dimension 3, hence they span R3.


ii)


"a\begin{pmatrix}
 1 \\
 1 \\
1
\end{pmatrix}+b\begin{pmatrix}
 0 \\
 2 \\
2
\end{pmatrix}+c\begin{pmatrix}
 0 \\
 0 \\
3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\
0
\end{pmatrix}"

"\begin{pmatrix}
 1 & 0 & 0 & & 0 \\
 1 & 2 & 0 & & 0 \\
 1 & 2 & 3 & & 0 \\
\end{pmatrix}\to\begin{pmatrix}
 1 & 0 & 0 & & 0 \\
 0 & 1 & 0 & & 0 \\
 0 & 0 & 1 & & 0 \\
\end{pmatrix}"

"a=b=c=0"

The given vectors are linearly independent.


iii) A subset "S" of a vector space "V"  is called a basis if

1. "S" is a spanning set

2. "S" is linearly independent.

Therefore the set "B=\{{(1,1,1),(0,2,2),(0,0,3)\}}" is a basis for R3.